Baldwin Company had 58,000 shares of common stock outstanding on January 1, 2021. On April 1, 2021, the company issued 38,000 shares of common stock. The company had outstanding fully vested incentive stock options for 28,000 shares exercisable at $10 that had not been exercised by its executives. The average market price of common stock for the year was $12. What number of shares of stock (rounded) should be used in computing diluted earnings per share?
Answer #1

Diluted EPS is is a profitability calculation which shows the amount of profit each share will receive if if all dilutive securities are executed.

It use weighted average number of commons shares and conversion of dilutive shares

(1) common shares

=existing shares+additional issue

=58,000+ 38,000*9/12 (as 38,000 shares are issued on April we will use 9 months)

=86,500 shares

(2) dilutive shares

=28,000*$10 /$12

=23,333

so, dilutive shares are 28,000-23,333 =4,667

Number of shares of stock (rounded) should be used in computing diluted earnings per share

=Common shares+dilutive shares

=86,500+4,667

=91,167